# Higher-Order Sinusoidal Input Describing Functions for Open-Loop and Closed-Loop Reset Control with Application to Mechatronics Systems

### v1

@@ 3.1 HOSIDFs for the Open-Loop Reset Control Systems
Consider an open-loop reset control system as shown in Fig. 4, with an input signal $e_{o}(t)=|E|\sin(\omega t+\angle E)$, resulting in the output signal $y_{o}(t)$ under Assumption 1. Using the “Virtual Harmonic Generator” [19], the input signal $e_{o}(t)$ generates harmonics expressed as $e_{o}^{n}(t)=|E|\sin(\omega t+n\angle E)$, with the corresponding Fourier transform denoted as $E_{o}^{n}(\omega)$. The signals $z_{o}(t)$, $m_{o}(t)$, and $y_{o}(t)$ consist of $n$ harmonics, represented as $z_{o}^{n}(t)$, $m_{o}^{n}(t)$, and $y_{o}^{n}(t)$, with Fourier transforms $Z_{o}^{n}(\omega)$, $M_{o}^{n}(\omega)$, and $Y_{o}^{n}(\omega)$, respectively. The Higher-Order Sinusoidal Input Describing Functions (HOSIDFs) of the reset controller $\mathcal{C}_{r}$ are given by
@@
$y_{o}(t)=\sum\nolimits_{n=1}^{\infty}y_{o}^{n}(t)=\sum\nolimits_{n=1}^{\infty}|E\mathcal{L}_{n}(\omega)|$ $\sin(n\omega t+\angle\mathcal{L}_{n}(\omega)),n=2k+1(k\in\mathbb{N}).$ (9)
@@ 4.1 HOSIDFs for the Closed-Loop Reset Control Systems
Theorem 2 and Corollary 1 establish the relationship between the HOSIDFs for open-loop and closed-loop reset control systems. The following remark illustrates this connection for the sensitivity functions $\mathcal{S}_{n}(\omega)$.
Remark 2.
The sensitivity functions $\mathcal{S}_{n}(\omega)$ in (11) for the closed-loop reset feedback control system in Fig. 1 are related to the open-loop transfer function $\mathcal{L}_{n}(\omega)$ in (7) and the base-linear sensitivity function $\mathcal{S}_{l}(\omega)$ in (14) as follows:
  1. 1.
    For the first-order harmonic ($n=1$), the sensitivity function is given by:
    $\mathcal{S}_{1}(\omega)=\frac{\mathcal{S}_{\text{DF}}(\omega)}{1+\Gamma_{\rho}(\omega)\mathcal{S}_{\text{DF}}(\omega)},$ (16)
    where
    $\mathcal{S}_{\text{DF}}(\omega)$ $=\frac{1}{1+\mathcal{L}_{1}(\omega)},$ (17)
    $\Gamma_{\rho}(\omega)$ $=(\Gamma(\omega)-1)\mathcal{L}_{\rho}(\omega).$
    Here, $\mathcal{S}_{\text{DF}}(\omega)$ represents the sensitivity function derived using the SIDF analysis [3]. From (16) and (17), when $\Gamma(\omega)=1$, it follows that $\mathcal{S}_{1}(\omega)=\mathcal{S}_{\text{DF}}(\omega)$. This indicates that the first-order harmonic $\mathcal{S}_{1}(\omega)$ is directly related to the first-order harmonic in the open-loop transfer function $\mathcal{L}_{1}(\omega)$, and the linear loop-shaping technique can be used to design the reset control systems [1].
  2. 2.
    For higher-order harmonics where $n>1$, the magnitude ratio to the first-order harmonic is expressed as:
    $\frac{|\mathcal{S}_{n}(\omega)|}{|\mathcal{S}_{1}(\omega)|}=\mathcal{S}_{l}(n\omega)\cdot\Gamma(\omega)\mathcal{L}_{n}(\omega)\mathcal{C}_{4}(n\omega).$ (18)
    From (18), for a reset control system with a settled base-linear sensitivity function $\mathcal{S}_{l}(n\omega)$, reducing $\Gamma(\omega)\mathcal{L}_{n}(\omega)\mathcal{C}_{4}(n\omega)$ in the open loop can decrease ${|\mathcal{S}_{n}(\omega)|}/{|\mathcal{S}_{1}(\omega)|}$ in the closed loop.
@@ 6.2 Discussions
The frequency response analysis methods in Theorems 1 to 2, developed for the generalized reset feedback control structure shown in Fig. 1, enable the tuning of linear elements in parallel, in series before and after the reset controller, and within the shaping filter to refine reset actions. This tuning flexibility broadens the potential for reset feedback control systems with enhanced performance. For example, as demonstrated in Remark 2, by ensuring that $\Gamma(\omega)\to 1$ in (14), the closed-loop sensitivity function can be approximately analyzed using classical linear SIDF and loop-shaping methods. Additionally, from (17), ensuring that ${|\mathcal{S}_{n}(\omega)|}/{|\mathcal{S}_{1}(\omega)|}\to 0$ can preserve the advantages of first-order harmonics while suppressing higher-order harmonics to negligible levels. Striking this balance enables the application of the superposition principle in closed-loop, multiple-input reset control systems.
@@ 7 Conclusion
Moreover, future research could leverage the frequency-domain techniques presented in Theorems 1 and 2, along with the insights provided in Remark 2, to develop guidelines for designing reset control systems. These guidelines could focus on optimizing both transient and steady-state responses in reset-controlled mechatronics systems.
@@ Appendix A The Proof of Theorem 1
Thus, from (37), (38), and (40), the $n$-th transfer function $\mathcal{L}_{n}(\omega)$ for the open-loop reset control system is given by
$\mathcal{L}_{n}(\omega)=\frac{Y_{o}^{n}(\omega)}{E_{o}^{n}(\omega)}=\begin{cases}\mathcal{C}_{1}(\omega)[\mathcal{C}_{r}^{1}(\omega)+\mathcal{C}_{2}(\omega)]\mathcal{C}_{3}(\omega)\mathcal{P}(\omega),&\text{for }n=1,\\ \mathcal{C}_{1}(\omega)e^{j(n-1)\angle\mathcal{C}_{1}(\omega)}\mathcal{C}_{r}^{n}(\omega)\mathcal{C}_{3}(n\omega)\mathcal{P}(n\omega),&\text{for odd }n>1,\\ 0,&\text{for even }n\geqslant 2.\end{cases}$ (41)
Finally, equations (31) and (41) complete the proof.
@@ Appendix B The Proof of Theorem 2
Consider a closed-loop reset control system in Fig. 1 with a sinusoidal reference input signal $r(t)=|R|\sin(\omega t)$, under Assumptions 2 and 3. This proof derives the HOSIDFs for this closed-loop system through three steps.
@@
$\Delta_{q}^{n}(\omega)$ $=(I+e^{A_{R}\pi/\omega})(A_{\rho}e^{A_{R}\pi/\omega}-I)^{-1}(I-A_{\rho})\Delta_{c}^{n}(\omega),$
@@
As shown in Fig. 17, the reset controller $\mathcal{C}_{r}$ processes distinct input signals $z^{n}(t)$ while relying on the same reset-triggered signal $z_{s}^{1}(t)$. According to (46), the nonlinear components $m^{n}_{\rho}(t)$ for different $n$ share identical phase and period. Then, define $v_{\rho}(t)=\sum\nolimits_{n=1}^{\infty}m^{n}_{\rho}(t)$ as the summation of nonlinear components $m^{n}_{\rho}(t)$. From (46), $V_{\rho}(\omega)=\mathscr{F}[v_{\rho}(t)]$ is given by:

### v3

@@ 3.1 HOSIDFs for the Open-Loop Reset Control Systems
Consider an open-loop reset control system as shown in Fig. 4, with an input signal $e_{o}(t)=|E|\sin(\omega t+\angle E)$, resulting in the output signal $y_{o}(t)$ under Assumption 1. Using the “Virtual Harmonic Generator” [19], the input signal $e_{o}(t)$ generates harmonics expressed as $e_{o}^{n}(t)=|E|\sin(n\omega t+n\angle E)$, with the corresponding Fourier transform denoted as $E_{o}^{n}(\omega)$. The signals $z_{o}(t)$, $m_{o}(t)$, and $y_{o}(t)$ consist of $n$ harmonics, represented as $z_{o}^{n}(t)$, $m_{o}^{n}(t)$, and $y_{o}^{n}(t)$, with Fourier transforms $Z_{o}^{n}(\omega)$, $M_{o}^{n}(\omega)$, and $Y_{o}^{n}(\omega)$, respectively. The Higher-Order Sinusoidal Input Describing Functions (HOSIDFs) of the reset controller $\mathcal{C}_{r}$ are given by
@@
$y_{o}(t)=\sum\nolimits_{n=1}^{\infty}y_{o}^{n}(t)=\sum\nolimits_{n=1}^{\infty}|E\mathcal{L}_{n}(\omega)|$ $\sin(n\omega t+n\angle E+\angle\mathcal{L}_{n}(\omega)),n=2k+1(k\in\mathbb{N}).$ (9)
@@ 6.2 Discussions
The frequency response analysis methods in Theorems 1 to 2, developed for the generalized reset feedback control structure shown in Fig. 1, enable the tuning of linear elements in parallel, in series before and after the reset controller, and within the shaping filter to refine reset actions. This tuning flexibility broadens the potential for reset feedback control systems with enhanced performance. For example, ensuring that ${|\mathcal{S}_{n}(\omega)|}/{|\mathcal{S}_{1}(\omega)|}\to 0$ can preserve the advantages of first-order harmonics while suppressing higher-order harmonics to negligible levels. Striking this balance enables the application of the superposition principle in closed-loop, multiple-input reset control systems.
@@ 7 Conclusion
Moreover, future research could leverage the frequency-domain techniques presented in Theorems 1 and 2 to develop guidelines for designing reset control systems. These guidelines could focus on optimizing both transient and steady-state responses in reset-controlled mechatronics systems.
CRediT authorship contribution statement
Xinxin Zhang: Conceptualization, Methodology, Formal analysis, Data curation, Software, Validation, Conducting Simulation and experiments, Writing – original draft, Writing – review and editing. S. Hassan HosseinNia: Conceptualization, Supervision, Resources, Review, and Editing.
@@
Acknowledgement
Xinxin Zhang acknowledges the PhD grant from the China Scholarship Council.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used [ChatGPT] in order to improve readability and language. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Appendix A The Proof of Theorem 1
The subsequent content builds upon the HOSIDF for the reset controller $\mathcal{C}_{r}$ in (28) to derive the HOSIDF in (7) for the open-loop reset system shown in Fig. 4. This derivation follows the same logical framework used in [22], but to accommodate the reset control structure presented in this study and ensure the completeness of the proof for (7), the following content is provided.
@@
Thus, from (28), (34), (35), and (37), the $n$-th transfer function $\mathcal{L}_{n}(\omega)$ for the open-loop reset control system is given by
$\mathcal{L}_{n}(\omega)=\frac{Y_{o}^{n}(\omega)}{E_{o}^{n}(\omega)}=\begin{cases}\mathcal{C}_{1}(\omega)[\mathcal{C}_{l}(\omega)+\mathcal{C}_{\rho}^{1}(\omega)+\mathcal{C}_{2}(\omega)]\mathcal{C}_{3}(\omega)\mathcal{P}(\omega),&\text{for }n=1,\\ \mathcal{C}_{1}(\omega)e^{j(n-1)\angle\mathcal{C}_{1}(\omega)}\mathcal{C}_{\rho}^{n}(\omega)\mathcal{C}_{3}(n\omega)\mathcal{P}(n\omega),&\text{for odd }n>1,\\ 0,&\text{for even }n\geqslant 2.\end{cases}$ (38)
Finally, equations (28) and (38) complete the proof. ∎
@@ Appendix B The Proof of Theorem 2
Consider a closed-loop reset control system, as shown in Fig. 1, with a sinusoidal reference input signal $r(t)=|R|\sin(\omega t)$, under Assumptions 2 and 3. This proof derives the HOSIDFs for the closed-loop system in three steps. The approach follows a similar process to the proof in [23], but while the latter is limited to a system with $\mathcal{C}_{1}=\mathcal{C}_{s}=\mathcal{C}_{3}=\mathcal{C}_{4}=1$ and $\mathcal{C}_{2}=0$, this proof is adapted to a more generalized structure as depicted in Fig. 1.
@@
$\Delta_{q}^{n}(\omega)$ $=(I+e^{A_{R}\pi/\omega})(A_{\rho}e^{A_{R}\pi/\omega}+I)^{-1}(I-A_{\rho})\Delta_{c}^{n}(\omega),$
@@
As shown in Fig. 17, the reset controller $\mathcal{C}_{r}$ processes distinct input signals $z^{n}(t)$ while relying on the same reset-triggered signal $z_{s}^{1}(t)$. Then, define $v_{\rho}(t)=\sum\nolimits_{n=1}^{\infty}m^{n}_{\rho}(t)$ as the summation of nonlinear components $m^{n}_{\rho}(t)$. From (43), $V_{\rho}(\omega)=\mathscr{F}[v_{\rho}(t)]$ is given by:
@@
According to (43), the nonlinear components $m^{n}_{\rho}(t)$ for different $n$ share identical phase and period. Therefore, $\Gamma(\omega)$ is a real number.